Write the slope-intercept equation for the line containing the given pair of numbers. (-6,1) (6,7)
Answer by stanbon(75887)

Write the slope-intercept equation for the line containing the given pair of numbers. (-6,1) (6,7)
---
slope = (7-1)/(6--6) = 6/12 = 1/2
------
Form:: y = mx + b
Solve for "b"::
1 = (1/2)(-6) + b
b = 1+3 = 4
------
Equation:
y = (1/2)x + 4
---------------
Cheers,
Stan H.
--------------